OpenStudy (anonymous):

The circle shown below has a diameter of 18 centimeters. What is the area of the shaded sector?

OpenStudy (mathstudent55):

\(\large A_{sector} = \dfrac{n}{360^o}\pi r^2\) where n = measure of sector, and r = radius of circle